Solve the following:
a. On a piece of paper, 8 points are marked such that no 3 points lie on the same straight line. How many lines can be drawn passing through any 2 of these points
b. Seven points lie on a circle. How many triangles can be drawn using any 3 of these points as vertices?
c. A group of 4 adults and 3 children are to be formed from 8 adults and 5 children. How many possible groups are there?
d. To promote reading, a teacher decides to feature 3 classics, 4 contemporary novels and 2 non-fiction books on a notice board. How many selections can she make from 5 classics, 6 contemporary novels and 4 non-fiction books?

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3817)

I'll do part (a) to get you started.

We have n = 8 points and we want to select r = 2 points from that list to generate a line.

There are nCr = 8C2 = 28 different ways to do this. The order of the two points selected doesn't matter. Segment AB is the same as segment BA.

--------------------

Here are the steps on how to compute 8C2 through the nCr combination formula
Before we do that though, let's note that
8! = 8*7*6*5*4*3*2*1
6! = 6*5*4*3*2*1
Both right hand sides contain "6*5*4*3*2*1" which allows us to say
8! = 8*7*6!

Now we can compute the nCr value
nCr = (n!)/(r!*(n-r)!)
8C2 = (8!)/(2!*(8-2)!)
8C2 = (8!)/(2!*6!)
8C2 = (8*7*6!)/(2!*6!) .... use what was discussed above
8C2 = (8*7)/(2!) .... the "6!" terms cancel
8C2 = 56/(2*1)
8C2 = 56/2
8C2 = 28

--------------------

Put another way:
We have 2 slots to fill and 8 items to pick from. The first slot has 8 choices and the second slot has 7 choices since we can't reuse whatever was picked in the first slot.
This gives 8*7 = 56 permutations where order matters (eg: AB would be different from BA)
However, naming a segment can be done in two different ways. So AB is not different from BA, and order doesn't matter. We have to divide by 2 to correct for the fact we double-counted
That's why we say 56/2 = 28.
There are 28 combinations possible.
So there are 28 different lines possible.

--------------------

Answer: 28

Answer by ikleyn(52781)

.

One and ONLY ONE problem/question per post.



It is the RULE, the POLICY and the REQUIREMENT of this forum.


It is written in this page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.


It is assumed that you read these rules before posting.

It is also assumed that you do understand what is written in that page and follow the rules.


Those who violate them, work against their own interests.